### v2

@@ Abstract
Simulating low-likelihood high-impact extreme weather events in a warming world is a significant and challenging task for current ensemble forecasting systems. While these systems presently use up to 100 members, larger ensembles could enrich the sampling of internal variability. They may capture the long tails associated with climate hazards better than traditional ensemble sizes. Due to computational constraints, it is infeasible to generate huge ensembles (comprised of 1,000-10,000 members) with traditional, physics-based numerical models. In this two-part paper, we replace traditional numerical simulations with machine learning (ML) to generate hindcasts of huge ensembles. In Part I, we construct an ensemble weather forecasting system based on Spherical Fourier Neural Operators (SFNO), and we discuss important design decisions for constructing such an ensemble. The ensemble represents model uncertainty through perturbed-parameter techniques, and it represents initial condition uncertainty through bred vectors, which sample the fastest growing modes of the forecast. Using the European Centre for Medium-Range Weather Forecasts Integrated Forecasting System (IFS) as a baseline, we develop an evaluation pipeline composed of mean, spectral, and extreme diagnostics. Using large-scale, distributed SFNOs with 1.1 billion learned parameters, we achieve calibrated probabilistic forecasts. As the trajectories of the individual members diverge, the ML ensemble mean spectra degrade with lead time, consistent with physical expectations. However, the individual ensemble members’ spectra stay constant with lead time. Therefore, these members simulate realistic weather states during the rollout, and the ML ensemble thus passes a crucial spectral test in the literature. The IFS and ML ensembles have similar Extreme Forecast Indices, and we show that the ML extreme weather forecasts are reliable and discriminating. hese diagnosticsy ensure that the ensemble can reliably simulate the time evolution of the atmosphere, including low likelihood high-impact extremes. In Part II, we generate a huge ensemble initialized each day in summer 2023, and we characterize the statistics of extremes.
@@ 1 Introduction
Recent low-likelihood, high-impact events (LLHIs) have raised important and unanswered questions about the drivers of these events and their relationship to anthropogenic climate change. For example, Hurricane Harvey in 2017 and the Summer 2021 heatwave in the Pacific Northwest (PNW) are two high-impact events with no modern analog. Several threads motivate research on LLHIs. First, the IPCC states that "the future occurrence of LLHI events linked to climate extremes is generally associated with low confidence" (Seneviratne et al., 2021, pp. 1536). Second, the occurrence of recent LLHIs, such as as the Summer 2021 PNW heatwave, reveals that the ability to characterize, let alone anticipate, such events is currently incomplete (Bercos-Hickey et al., 2022; Zhang et al., 2024; Liu et al., 2024).
@@
These challenges motivate the application of entirely new methodological approaches, such as those based on machine learning (ML). For the first time, it is now possible to generate massive ensembles using ML with orders-of-magnitude less computational cost than traditional numerical simulations (Pathak et al., 2022). Recent work has demonstrated the potential of deterministic ML-based weather forecasting, which has comparable or superior root-mean squared error (RMSE) to the Integrated Forecasting System (IFS) at 0.25 degree horizontal resolution (Bi et al., 2023; Lam et al., 2023; Willard et al., 2024; ECMWF, ). As our ML architecture, we use Spherical Fourier Neural Operators (SFNO) (Bonev et al., 2023). SFNO has been proven to be efficient and powerful in modeling a wide range of chaotic dynamical systems, including turbulent flows and atmospheric dynamics, while remaining numerically stable over long autoregressive rollouts. Given these promising deterministic results, we use ML to create ensemble forecasts, which provide probabilistic weather predictions. A high-level design decision is whether to create the ensemble after training the ML model or during the training itself. We use the former approach: we train ML models to minimize the deterministic mean squared error (MSE) at each time step. After training, we create a calibrated ensemble by representing initial condition and model uncertainty. Conversely, NeuralGCM (Kochkov et al., 2023), FuXi-ENS (Zhong et al., 2024), SEEDS (Li et al., 2024), and GenCast (Price et al., 2023) employ probabilistic training objectives instead of deterministic RMSE.
In this two-part paper, we present a first-of-its-kind huge ensemble of weather extremes using an ML-based emulator of global numerical reanalyses. In Part I, we introduce the ML architecture and the ensemble design (Section 2). In Table 1, we list the major design decisions of the ensemble, and we include pointers to the relevant sections in the paper for understanding the decision-making criteria. We benchmark the ML performance against an operational weather forecast, the European Center for Medium-range Weather Forecast’s (ECMWF) Integrated Forecasting System ensemble (IFS ENS). We assess whether our ML ensemble is fit for purpose using a suite of diagnostics that assess the overall probabilistic performance of the ensemble and its spectra. Because of our interest in LLHIs, we also present an extremes diagnostics pipeline that specifically assesses ML extreme weather forecasts. In Part II, we analyze a huge ensemble with hindcasts of 7,424 ensemble members.
@@ 2 Designing ensembles with SFNO
We adopt the SFNO training scheme presented in Bonev et al. (2023). SFNO is trained on the European Center for Medium-range Weather Forecasts Reanalysis v5 (ERA5) (Hersbach et al., 2020) at the dataset’s 0.25-degree horizontal resolution. The weights of SFNO are optimized to minimize the latitude-weighted deterministic MSE loss function. Each model is trained for 70 epochs
@@
To create ensemble weather forecasts with SFNO, we mirror methods used in numerical weather prediction. Figure 1 provides an overview of our ensemble generation method. For weather forecasts, two major sources of uncertainty are initial condition uncertainty and model uncertainty. Initial condition uncertainty stems from the inaccuracies in observing the current meteorological state, while model uncertainty arises from the incompletely known and imperfect numerical representations of physics that govern the atmosphere’s time evolution. To represent initial condition uncertainty, we use bred vectors, a method formerly used by the Global Ensemble Forecast System (GEFS) (Toth and Kalnay, 1993, 1997). Bred vectors are designed to sample the fastest growing directions of the ensemble error patterns. By creating rapidly diverging ensemble trajectories, bred vectors are designed to create an ensemble fully represents the probability of future weather states. Existing work has shown that simple Gaussian perturbations do not yield a sufficiently dispersive ensemble (Scher and Messori, 2021; Bülte et al., 2024): the ensemble spread from these perturbations is too small. Bred vectors solve this problem by creating a more dispersive ensemble which better reflects the full distribution of possible future states of the atmosphere. They agnostically amplify the fastest growing modes in a model’s intrinsic dynamics. While bred vectors have been used to create ensemble forecasts from traditional dynamical models, assessing how ML models respond to perturbations is an important research frontier.
@@
To represent model uncertainty, we train multiple SFNO models from scratch. We refer to each trained SFNO instance as a "checkpoint." At the start of training, each checkpoint is initialized with different random weights. During training, SFNO iteratively updates its weights to minimize a loss function: in this case, the loss function is the mean-squared error between the model predictions and the ERA5 training data. During each epoch of training, SFNO iterates through the entire training dataset and updates its weights to minimize the loss. We train SFNO for 70 total epochs. By the end of training, the models converge to a different local optimum of learned weights. The resulting ensemble of the different trained SFNO checkpoints represents the uncertainty in the SFNO model weights itself. Each resulting checkpoint represents an equivalently plausible set of weights that can model the time evolution of the atmosphere from an initial state. With multiple checkpoints, we create an ensemble with a spread of forecasts, yet each ensemble member has the same skill. Weyn et al. (2021) use multiple checkpoints to create an ensemble of forecasting models for medium-range and subseasonal prediction. They reduce computational costs by saving multiple model checkpoints from each training run and training the last few epochs independently for each model. This approach requires several additional design decisions: how should the learning rate for the optimization during these last retrained epochs be adjusted? How many extra epochs should each checkpoint train for? At what point during training should the checkpoints diverge? To minimize the ensemble’s dependence on these hyperparameters, we opt to retrain each checkpoint completely from scratch.
@@
We create an ensemble called SFNO-BVMC: Spherical Fourier Neural Operators with Bred Vectors and Multiple Checkpoints. In Table 1, we present a list of hyperparameters and their associated criteria that we use to guide our choice of ensemble design. We use a train-validation-test set paradigm. SFNO is trained on the years 1979-2016. We use the year 2018 as a validation year, on which we tune multiple aspects of the ensemble, such as the amplitude of the bred vectors and the number of SFNO checkpoints. Because these ensemble parameters are tuned using the year 2018, we cannot use 2018 for unbiased evaluation of the final ensemble. For our overall diagnostics, the year 2020 is used as an out-of-sample, held-out test set This year is used in the test set on the WeatherBench 2 (Rasp et al., 2024) platform, allowing for simplified comparison of SFNO-BVMC with other ML-based ensemble weather forecasting systems. To evaluate the skill for extreme weather, we use boreal summer 2023 (June, July, August) because it is the hottest summer in recorded history (Esper et al., 2024). In Part II, we present a deep dive on a huge ensemble of forecasts from this time period. No SFNO training or ensemble design decisions were made using the year 2020 or 2023. This setup with different training, validation, and test sets is crucial to avoid data leakage.
@@
Number of Perturbations per Model Checkpoint (Benchmark Ensemble) 2, with 1 centered bred vector perturbation Part I, Section 3
Number of perturbations per Model Checkpoint (Huge Ensemble) 256, with 128 centered bred vector perturbations Part II
@@
Lead Time to Analyze Extreme Statistics 3–5 days, 10 days Part I, Section 3.3
@@ 2.1 Selecting an emulator
SFNO is an ML architecture built on neural operators (Li et al., 2020), which are designed to learn mappings between function spaces. They can be used for different discretizations and grids and have broad applicability to various partial differential equation (PDE) problems. SFNO is a special instance of a Neural Operator, which uses the Spherical Harmonic Transform to represent operators acting on functions defined on the sphere. The spherical formulation leads to a strong inductive bias, respecting underlying symmetries of the problem. This reduces error buildup and leads to stable autoregressive rollouts, making the methods ideally suited for PDE problems on the sphere. We use the open-source version of SFNO v0.1.0 released in modulus-makani (Bonev et al., 2024). The input to SFNO consists of seventy-four channels comprising the meteorological state at a given time. The model then predicts those same seventy-four channels at a future time of six hours, which also determines the time step of the SFNO-BVMC ensemble. In addition to the prognostic channels, we add three extra input channels: the cosine of the solar zenith angle, orography, and land-sea mask.
@@
The existing implementation of SFNO from Bonev et al. (2023) makes forecasts for seventy-three total prognostic variables. In this study, we add 2-meter (2m) dewpoint temperature as another variable; for our SFNO training dataset, we obtain the 2m dewpoint temperature field from ERA5. Together, 2m dewpoint temperature and 2m air temperature provide an estimate of heat and humidity at the surface. Since we have trained SFNOs to predict both these variables, we can assess LLHI heat-humidity events. It is vital to assess the combination of both heat and humidity to characterize heat stress and LLHIs in a warming world (Vargas Zeppetello et al., 2022). Olivetti and Messori (2024) evaluate deterministic ML-based extreme weather forecasts, but they use 1000 hPa specific humidity as a proxy for surface humidity. They note that this approximation has limitations in predicting the surface heat stress and heat index. We build on their work by creating a model that predicts 2m temperature and dewpoint; also, while they analyze deterministic forecasts, our diagnostics pipeline is designed for extreme weather forecasts from ensembles. The addition of 2m dewpoint also enables estimating Convective Available Potential Energy in the forecasts from SFNO. By quantifying the buildup of convective instability, this variable is useful for studying convective storms and thunderstorms. In total, we train SFNO to predict 74 meteorological variables, which are listed in Table 2.
@@
We choose SFNO because its spherical design is well-suited for problems in earth science, and the architecture includes scalable model parallel implementations (Bonev et al., 2023; Kurth et al., 2023). Since SFNO can be split across multiple GPUs during training, we can train large SFNOs and assess the effect of the SFNO size on the ensemble dispersion.
SFNO contains a number of hyperparameters that determine the total size of the model and its ensemble performance. Two such hyperparameters are the scale factor and the embedding dimension. The scale factor controls the level of spectral downsampling of the input field. With larger downsampling, the effective resolution of the SFNO decreases, and finer scale perturbations are blurred out. These perturbations would not grow appreciably during the model rollout, so the model spread would not cover the range of future weather states. Thus, we expect a model with a lower scale factor (less downsampling) to have larger ensemble spread. The embedding dimension determines the size of the learned representation of the input fields (Pathak et al., 2022). A larger embedding dimension increases the number of learnable parameters in the SFNO, thereby requiring more GPU memory.
@@
Figure 2: Comparing different versions of SFNO. (a) The 850 hPa temperature spread-error ratios are compared for lagged ensembles. A lagged ensemble is created by initializing SFNO using nine adjacent time steps, and the spread-error is shown for each SFNO configurations. (b) Relative power spectra (colored lines) for 850 hPa temperatures for a large SFNO (with a scale factor of 2 and an embed dimension of 620), a medium-sized SFNO (scale factor 4 and embed dimension 384), and a small SFNO (scale factor 6 and embed dimension 220). Spectra are computed relative to the ERA5 spectrum (horizontal black line).
@@
To select an SFNO architecture, we assess how these hyperparameters affect lagged ensemble spread-error ratio and spectral degradation. A lagged ensemble creates an ensemble by using nine adjacent time steps as initial conditions (Brankovic et al., 1990). Brenowitz et al. (2024) analyze lagged ensembles of deterministic ML weather models to assess their intrinsic dispersiveness. A spread-error ratio can be calculated from this ensemble. Ordinarily, benchmarking the ensemble performance would require tuning the ensemble parameters (e.g. amplitude of perturbations, number of checkpoints, form of perturbations) separately for each architecture. This process is time-consuming, memory-intensive, and computationally demanding. Lagged ensembles enable comparison of different deterministic architectures without separately tuning ensembling methods for each architecture. In Figure 2a, the lagged ensemble spread-error ratio for 850 hPa temperature is highest for the large model, indicating that this model is best-suited for ensemble forecasting. The spread-error ratio systematically improves for the larger models. Brenowitz et al. (2024) find complementary results; they show that smaller scale factors favorably enhance dispersion. Here, we consider the combined effect of changing both scale factor and embedding dimension.
@@
We compare the small, medium, and large SFNOs’ spectra, and we assess the extent to which they fully resolve the spectrum of the underlying ERA5 data. A known problem with deterministic ML models is that the small wavelengths are blurry (Kochkov et al., 2023). We attempt to suppress this blurring as much as possible by using a small scale factor and a large embedding dimension. By design, a small scale factor and large embed dimension allow the model’s latent state to represent the input with more parameters and higher resolution. Figure 2b shows that the larger models (with lower scale factors and larger embed dimension) have less spectral degradation and are better able to preserve the spectra of ERA5. Based on these two tests, we use the large version of SFNO, with a scale factor of 2 and embed dimension of 620. This version of SFNO trains in 16 hours on 256 80GB NVIDIA A100 GPUs. It leverages data parallelism, in which the batch size of 64 is split up across different GPUs, and model parallelism, in which the input field is divided into four latitude bands. These four bands are split across four GPUs (one GPU per section), and the SFNO architecture is distributed to train with spatial model parallelism. Each SFNO checkpoint trains for 70 epochs using a pressure-weighted mean squared error loss function (Lam et al., 2023).
@@
We note that there are many possible combinations of the scale factor, embedding dimension, and other training hyperparameters. We do not conduct comprehensive hyperparameter tuning via a grid search. Such an experiment would be very computationally expensive, due to the large number of hyperparameter combinations. Instead, we optimize the scale factor and embedding dimension because of their direct relevance to spectral degradation. Instead of hyperparameter tuning, we choose to expend our compute budget on training as many checkpoints as possible. With multiple checkpoints, we aim to span the model space of all possible SFNO checkpoints that have our chosen architecture and hyperparameters. By having as many SFNO checkpoints as possible, we hope to increase our coverage of extreme weather events with a thorough representation of model uncertainty.

### v3

@@ Abstract
Simulating low-likelihood high-impact extreme weather events in a warming world is a significant and challenging task for current ensemble forecasting systems. While these systems presently use up to 100 members, larger ensembles could enrich the sampling of internal variability. They may capture the long tails associated with climate hazards better than traditional ensemble sizes. Due to computational constraints, it is infeasible to generate huge ensembles (comprised of 1,000-10,000 members) with traditional, physics-based numerical models. In this two-part paper, we replace traditional numerical simulations with machine learning (ML) to generate hindcasts of huge ensembles. In Part I, we construct an ensemble weather forecasting system based on Spherical Fourier Neural Operators (SFNO), and we discuss important design decisions for constructing such an ensemble. The ensemble represents model uncertainty through perturbed-parameter techniques, and it represents initial condition uncertainty through bred vectors, which sample the fastest growing modes of the forecast. Using the European Centre for Medium-Range Weather Forecasts Integrated Forecasting System (IFS) as a baseline, we develop an evaluation pipeline composed of mean, spectral, and extreme diagnostics. With large-scale, distributed SFNOs with 1.1 billion learned parameters, we achieve calibrated probabilistic forecasts. As the trajectories of the individual members diverge, the ML ensemble mean spectra degrade with lead time, consistent with physical expectations. However, the individual ensemble members’ spectra stay constant with lead time. Therefore, these members simulate realistic weather states during the rollout, and the ML ensemble passes a crucial spectral test in the literature. The IFS and ML ensembles have similar Extreme Forecast Indices, and we show that the ML extreme weather forecasts are reliable and discriminating. These diagnostics ensure that the ensemble can reliably simulate the time evolution of the atmosphere, including low likelihood high-impact extremes. In Part II, we generate a huge ensemble initialized each day in summer 2023, and we characterize the simulations of extremes.
@@ 1 Introduction
Recent low-likelihood, high-impact events (LLHIs) have raised important and unanswered questions about the drivers of these events and their relationship to anthropogenic climate change. For example, Hurricane Harvey in 2017 and the Summer 2021 heatwave in the Pacific Northwest (PNW) are two high-impact events with no modern analog. Several threads motivate research on LLHIs. First, the IPCC states that "the future occurrence of LLHI events linked to climate extremes is generally associated with low confidence" (Seneviratne et al., 2021, pp. 1536). Second, the occurrence of recent LLHIs, such as the Summer 2021 PNW heatwave, reveals that our abilities to characterize, let alone anticipate, such events are currently incomplete (Bercos-Hickey et al., 2022; Zhang et al., 2024; Liu et al., 2024).
@@
These challenges motivate the application of entirely new methodological approaches, such as those based on machine learning (ML). For the first time, it is now possible to generate massive ensembles using ML with orders-of-magnitude less computational cost than traditional numerical simulations (Pathak et al., 2022). Recent work has demonstrated the potential of deterministic ML-based weather forecasting, which has comparable or superior root-mean squared error (RMSE) to the Integrated Forecasting System (IFS) at 0.25 degree horizontal resolution (Bi et al., 2023; Lam et al., 2023; Willard et al., 2024; ECMWF, ). Olivetti and Messori (2024) show that these deterministic data-driven models also offer promising forecast skill on extremes, and Pasche et al. (2024) validate them on case studies of extreme events. As our ML architecture, we use Spherical Fourier Neural Operators (SFNO) (Bonev et al., 2023). SFNO has been proven to be efficient and powerful in modeling a wide range of chaotic dynamical systems, including turbulent flows and atmospheric dynamics, while remaining numerically stable over long autoregressive rollouts. Given these promising deterministic results, we use ML to create ensemble forecasts, which provide probabilistic weather predictions. A high-level design decision is whether to create the ensemble after training the ML model or during the training itself. We use the former approach: we train ML models to minimize the deterministic mean squared error (MSE) at each time step. After training, we create a calibrated ensemble by representing initial condition and model uncertainty. Conversely, NeuralGCM (Kochkov et al., 2023), FuXi-ENS (Zhong et al., 2024), SEEDS (Li et al., 2024), and GenCast (Price et al., 2023) employ probabilistic training objectives instead of deterministic RMSE.
In this two-part paper, we present a first-of-its-kind huge ensemble of weather extremes using an ML-based emulator of global numerical reanalyses. In Part I, we introduce the ML architecture and the ensemble design (Section 2). In Table 1, we list the major design decisions of the ensemble, and we include pointers to the relevant sections in the paper for understanding the decision-making criteria. We benchmark the ML performance against an operational weather forecast, the European Center for Medium-range Weather Forecast’s (ECMWF) Integrated Forecasting System ensemble (IFS ENS). We assess whether our ML ensemble is fit for purpose using a suite of diagnostics that assess the overall probabilistic performance of the ensemble and its spectra. Because of our interest in LLHIs, we also present an extremes diagnostics pipeline that specifically assesses ensemble extreme weather forecasts. In Part II (Mahesh et al., 2024), we analyze a huge ensemble hindcast with 7,424 ensemble members.
@@ 2 Designing ensembles with SFNO
We adopt the SFNO training scheme presented in Bonev et al. (2023). SFNO is trained on the European Center for Medium-range Weather Forecasts Reanalysis v5 (ERA5) (Hersbach et al., 2020) at the dataset’s 0.25-degree horizontal resolution. The weights of SFNO are optimized to minimize the latitude-weighted deterministic MSE loss function. When calculating the loss, each variable is weighted by pressure and by the time tendency of the variable in the training dataset, similar to methods presented in Lam et al. (2023) and Watt-Meyer et al. (2023).
@@
To create ensemble weather forecasts with SFNO, we mirror methods used in numerical weather prediction. Figure 1 provides an overview of our ensemble generation method. For weather forecasts, two major sources of uncertainty are initial condition uncertainty and model uncertainty. Initial condition uncertainty stems from the inaccuracies in observing the current meteorological state, while model uncertainty arises from the incompletely known and imperfect numerical representations of physics that govern the atmosphere’s time evolution. To represent initial condition uncertainty, we use bred vectors, a method formerly used by the Global Ensemble Forecast System (GEFS) (Toth and Kalnay, 1993, 1997). Bred vectors are designed to sample the fastest growing directions of the ensemble error patterns. By creating rapidly diverging ensemble trajectories, bred vectors are designed to create an ensemble that fully represents the probability of future weather states. Existing work has shown that simple Gaussian perturbations do not yield sufficiently dispersive ML ensembles (Scher and Messori, 2021; Bülte et al., 2024): the ensemble spread from these perturbations is too small. Bred vectors help address this problem by creating a more dispersive ensemble which better reflects the full distribution of possible future states of the atmosphere. They amplify the fastest growing modes in the model’s intrinsic dynamics. While bred vectors have been used to create ensemble forecasts from traditional dynamical models, assessing how ML models respond to such perturbations is an important research frontier.
@@
To represent model uncertainty, we train multiple SFNO models from scratch. We refer to each trained SFNO instance as a "checkpoint." At the start of training, each checkpoint is initialized with different random weights. During training, SFNO iteratively updates its weights to minimize a loss function: in this case, the loss function is the mean-squared error between the model predictions and the ERA5 training data. During each epoch of training, SFNO iterates through the entire training dataset and updates its weights to minimize the loss. We train SFNO for 70 total epochs. By the end of training, the models converge to a different local optimum of learned weights. The resulting ensemble of the different trained SFNO checkpoints represents the uncertainty in the SFNO model weights itself. Each resulting checkpoint represents an equivalently plausible set of weights that can model the time evolution of the atmosphere from an initial state. With multiple checkpoints, we create an ensemble with a spread of forecasts. Weyn et al. (2021) use multiple checkpoints to create an ensemble of forecasting models for medium-range and subseasonal prediction. They reduce computational costs by saving multiple checkpoints from each training run and training the last few epochs independently for each model. This approach requires several additional design decisions: how should the learning rate for the optimization during these last retrained epochs be adjusted? How many extra epochs should each checkpoint train for? At what point during training should the checkpoints diverge? To minimize the ensemble’s dependence on these hyperparameters, we opt to retrain each checkpoint completely from scratch.
@@
We create an ensemble called SFNO-BVMC: Spherical Fourier Neural Operators with Bred Vectors and Multiple Checkpoints. In Table 1, we present a list of hyperparameters and their associated criteria that we use to guide our choice of ensemble design. We use a train-validation-test set paradigm. SFNO is trained on the years 1979-2016. We use the year 2018 as a validation year, on which we tune multiple aspects of the ensemble, such as the amplitude of the bred vectors and the number of SFNO checkpoints. Because these ensemble parameters are tuned using the year 2018, we cannot use 2018 for unbiased evaluation of the final ensemble. For our overall diagnostics, the year 2020 is used as an out-of-sample, held-out test set. To evaluate the skill for extreme weather, we use boreal summer 2023 (June, July, August) because it was the hottest summer in recorded history at the time (Esper et al., 2024). In Part II, we present a deep dive on a huge ensemble of forecasts from this time period. No SFNO training or ensemble design decisions were made using the year 2020 or 2023. This setup with different training, validation, and test sets is crucial to avoid data leakage.
@@
Number of Perturbations per Model Checkpoint (Benchmark Ensemble) 2 (1 bred vector perturbation that is added to and subtracted from the initial condition) Part I, Section 3
Number of perturbations per Model Checkpoint (Huge Ensemble) 256 (128 bred vector perturbations, each added to and subtracted from the initial condition) Part II
@@
Lead Time to Analyze Extreme Statistics 2 days, 4 days, 10 days Part I, Section 3.3
@@ 2.1 Selecting an emulator
SFNO is an ML architecture built on neural operators (Li et al., 2020), which are designed to learn mappings between function spaces. They can be used for different discretizations and grids, and they have broad applicability to various partial differential equation (PDE) problems. SFNO is a special instance of a Neural Operator, which uses the Spherical Harmonic Transform to represent operators acting on functions defined on the sphere. The spherical formulation leads to a strong inductive bias, respecting the geometry and symmetry of the sphere. This reduces error buildup during autoregressive rollouts. We use the open-source version of SFNO v0.1.0 released in the modulus-makani Python repository (Bonev et al., 2024).
We provide a brief overview of the SFNO architecture; for a more detailed explanation, refer to Bonev et al. (2023). The SFNO architecture consists of three main components: the encoder, the SFNO blocks, and the decoder.
  1. 1.
    Encoder: The encoder employs multi-layer perceptrons (MLPs) at each grid cell to map the input fields into a higher-dimensional latent space. MLPs are fully connected neural networks that apply nonlinear transformations to their inputs.
  2. 2.
    SFNO Blocks: The processor incorporates 8 SFNO blocks, operating in the latent space, each performing two main operations:
    1. (a)
      A spherical convolution with a learned filter encoded in the spectral domain. The signal is transformed into the spectral domain and back via a spherical harmonic transform (SHT) and its inverse. In the spectral domain, the convolution operation becomes a pointwise multiplication.
    2. (b)
      An MLP applies nonlinear transformations to the latent features.
    The output of each SFNO block serves as the input to the subsequent block. The first block downsamples the input resolution by a specified "scale factor," while the last block upsamples back to the original resolution.
  3. 3.
    Decoder: The decoder maps the latent space back into physical space using MLPs.
SFNO encodes an operator that maps functions defined on the sphere to other functions on the sphere. This learned map is parameterized by the weights of the MLPs, spectral filters, encoder, and decoder. These weights of SFNO are optimized during training.
The input to SFNO consists of seventy-four channels comprising the meteorological state at a given time (Table 2). The model then predicts those same seventy-four channels six hours in the future. In addition to the prognostic channels, we add three extra input channels: the cosine of the solar zenith angle, orography, and land-sea mask.
@@
The existing implementation of SFNO from Bonev et al. (2023) makes forecasts for seventy-three total prognostic channels. In this study, we add ERA5 2-meter (2m) dewpoint temperature as another variable. Together, 2m dewpoint temperature and 2m air temperature provide an estimate of heat and humidity at the surface. Since we have trained SFNOs to predict both these variables, we can simulate LLHI heat-humidity events. It is vital to assess the combination of both heat and humidity to characterize heat stress and LLHIs in a warming world (Vargas Zeppetello et al., 2022). While some ML models have 1000-hPa specific humidity, Pasche et al. (2024) note that this approximation has limitations in predicting the surface heat stress and heat index. Therefore, we add 2m dewpoint to more directly characterize moisture near the surface. In future work, the addition of 2m dewpoint could enable estimating Convective Available Potential Energy in the forecasts from SFNO. By quantifying the buildup of convective instability, this variable is useful for studying convective storms and thunderstorms.
@@
The SFNO architecture includes scalable model parallel implementations, in which the model is split across multiple GPUs during training (Bonev et al., 2023; Kurth et al., 2023). We train large SFNOs and assess the effect of the SFNO size on ensemble dispersion. SFNO contains a number of hyperparameters that determine the total size of the model and its ensemble performance. Two such hyperparameters are the scale factor and the embedding dimension. The scale factor specifies how much the input field is spectrally downsampled when creating the latent representation. With more aggressive downsampling, SFNO internally represents the input atmospheric state with reduced resolution. We speculate that this may reduce the effective resolution of the predictions (Brenowitz et al., 2024). With a reduced effective resolution, small-scale perturbations would not grow and propagate upscale. Instead, they would be blurred out, and they would not result in increased spread among ensemble members. The embedding dimension determines the size of the learned representation of the input fields (Pathak et al., 2022). A larger embedding dimension increases the number of learnable parameters in the SFNO, thereby requiring more GPU memory.
@@
Figure 2: Comparing different versions of SFNO. (a) The 850 hPa temperature spread-error ratios are compared for lagged ensembles. A lagged ensemble is created by using nine adjacent time steps as initial conditions, and the spread-error is shown for each SFNO configurations. (b) Relative power spectra at a lead time of 360 hours (colored lines) for 850 hPa temperatures for a large SFNO (with a scale factor of 2 and an embed dimension of 620), a medium-sized SFNO (scale factor 4 and embed dimension 384), and a small SFNO (scale factor 6 and embed dimension 220). Spectra are computed relative to the ERA5 spectrum (horizontal black line).
@@
To select an SFNO architecture, we assess how these hyperparameters affect lagged ensemble spread-error ratio and spectral degradation. A lagged ensemble is created by using nine adjacent time steps as initial conditions (Brankovic et al., 1990). Brenowitz et al. (2024) analyze the spread-error ratio of lagged ensembles to assess the intrinsic dispersion of deterministic ML models. Ordinarily, benchmarking ensemble performance would require generating and tuning a full set of ensemble parameters (e.g. amplitude of perturbations, number of checkpoints, form of perturbations) separately for each architecture. This process is time-consuming, memory-intensive, and computationally demanding. Lagged ensembles readily enable comparison of different deterministic architectures with minimal tuning parameters. In Figure 2a, the lagged ensemble spread-error ratio for 850 hPa temperature is closest to 1 for the large model, indicating that this model is best-suited for ensemble forecasting. The spread-error ratio systematically improves for the larger models. Brenowitz et al. (2024) find complementary results; they show that smaller scale factors improve the spread-error ratio. Here, we consider the combined effect of changing both scale factor and embedding dimension.
We assess the extent to which the small, medium, and large SFNOs fully resolve the spectrum of the ERA5 training data. A known problem with deterministic ML models is that the small wavelengths are blurry (Kochkov et al., 2023). We attempt to suppress this blurring as much as possible by using a small scale factor and a large embedding dimension. In addition, we intentionally avoid using autoregressive training (Lam et al., 2023; Pathak et al., 2022; Keisler, 2022). In this method (sometimes called "multistep finetuning" or "multistep loss"), the ML model weights are optimized over multiple timesteps, not just a 1-step prediction. The goal of this method is to improve the forecast performance by training the ML model to perform well when autoregressively rolled out with its own predictions. Brenowitz et al. (2024) and Lang et al. (2024) hypothesize that autoregressive training could contribute to spectral degradation. This method may effectively increase the time step of the model, making it more similar to an ensemble mean (Lang et al., 2024). Many deterministic models’ initial 1-step forecasts are blurry, and with this method, their forecasts get increasingly blurry with lead time. Because we do not use autoregressive finetuning, we hypothesize that SFNO has spectra that stay constant with lead time (see Section 3.2 for more discussion).
With this design decision, we also reduce the computational requirements of training SFNO. Autoregressive fine-tuning requires significant GPU memory and computation time because it calculates gradients across multiple model steps. With these computational savings, we train large SFNOs with a small scale factor and a large embedding dimension. These design choices allow our configuration of SFNO to hold as much high-resolution information in its internal representation as possible.
@@
Figure 2b shows that the larger models (with lower scale factors and larger embed dimension) have less spectral degradation and are better able to preserve the spectra of ERA5. Based on Figure 2a and b, we select the large version of SFNO as our final set of hyperparameters. This version of SFNO trains in 16 hours on 256 80GB NVIDIA A100 GPUs. It leverages data parallelism, in which the batch size of 64 is split up across different GPUs, and spatial model parallelism, in which the input field is divided into four latitude bands. Each SFNO checkpoint trains for 70 epochs using a pressure-weighted mean squared error loss function (Lam et al., 2023).
@@
We note that there are many possible combinations of the scale factor, embedding dimension, and other training hyperparameters. We do not conduct comprehensive hyperparameter tuning via a grid search. Such an experiment would be very computationally expensive, due to the large number of combinations. Instead, we optimize the scale factor and embedding dimension because of their direct relevance to spectral degradation. Rather than hyperparameter tuning, we choose to expend our compute budget on training as many checkpoints as possible to try to span the space of all possible SFNO checkpoints. With many SFNO checkpoints, we hope to increase our coverage of extreme weather events with a thorough representation of model uncertainty.